OpenStudy (anonymous):

write the equation of the line through (4,-7) and (2,0)

OpenStudy (anonymous):

y=ax+b a=delta y / delta x = (0--7)/(2-4) = 7/-2 = -3.5 so y=-3.5x +b we fill in one point: 0=-3.5 * 2 + b 0 = -7 + b so b=7 y=-3.5x + 7